# Lrq-Fact: LLM-Generated Relevant Questions for Multimodal Fact-Checking

### v1

@@ Abstract
Human fact-checkers have specialized domain knowledge that allows them to formulate precise questions to verify information accuracy. However, this expert-driven approach is labor-intensive and is not scalable, especially when dealing with complex multimodal misinformation. In this paper, we propose a fully-automated framework, Lrq-Fact, for multimodal fact-checking. First, the framework leverages Vision-Language Models (VLMs) and Large Language Models (LLMs) to generate comprehensive questions and answers for probing multimodal content. Next, a rule-based decision-maker module evaluates both the original content and the generated questions and answers to assess the overall veracity. Extensive experiments on two benchmarks show that Lrq-Fact achieves promising performance for multimodal misinformation detetcion. Moreover, we evaluate its generalizability across different model backbones, offering valuable insights for further refinement.
@@ 1 Introduction
Fact-checkers are experts with sufficient domain knowledge in their respective fields, enabling them to ask relevant and targeted questions to verify the factuality of information [72, 73]. This expertise allows them to assess the veracity of claims from multiple perspectives by formulating precise questions that probe both textual and visual content [55]. However, this manual approach is time-consuming, labor-intensive, and lacks scalability, especially given the increasing complexity and volume of multimodal misinformation. As misinformation continues to evolve, relying solely on manual efforts is no longer practical [8, 25, 27].
Figure 1: The proposed framework, Lrq-Fact, draws on insights from manual fact-checking process.
In response to this challenge, numerous automatic misinformation detection systems have recently emerged to address the problem [54]. While these systems provide valuable insights, most of them focus on textual misinformation only [81, 29, 8, 62]. This text-only detection approach may overlook critical visual cues. Few works [86, 22, 59] for multimodal scenarios often analyze these modalities in isolation, and then combine predictions from text and images data, thus leading to overlooking of fine-grained inconsistencies or manipulation patterns, such as minor facial expressions [37].
To overcome these limitations, we propose LLM-generated Relevant Questions for multimodal FACT-checking (Lrq-Fact), a framework designed to jointly consider the intricate inter-connections between textual and visual content. Lrq-Fact is inspired by the process of human fact-checkers, which can be abstracted into investigating two types of questions: those focused on image content and those focused on textual content [75]. Image-focused questions enable the system to carefully examine the visual content within the context of the article, ensuring that the image accurately reflects critical details such as people, objects, or events mentioned in the text. This step is crucial, as visual elements can be misleading or manipulated, even when the accompanying text appears accurate. Meanwhile, text-specific questions are used to verify the factual accuracy and plausibility of the claims made in the article itself.
Lrq-Fact incorporates the power of existing Vision-Language Models (VLMs) and Large Language Models (LLMs) to make the process of asking fact-verifying questions more scalable and efficient. LLMs generate precise and contextually relevant questions aimed at scrutinizing the textual content, while VLMs analyze visual content to answer image-focused queries. By integrating these models in a unified framework, Lrq-Fact addresses the limitations of isolated text or image analysis, enabling deeper detection of inconsistencies across modalities. In addition to generating separate questions and answers for image and text, our framework uses a rule-based decision-maker, where LLMs examine the insights from both modalities to determine the overall veracity of the content. This approach automates the fact-checking process and provides explainable rationales for the decisions, enhancing transparency and trust in the detection outcomes.
We evaluate the performance of our framework through extensive experiments on two benchmarks for multimodal misinformation detection. The results show that our method significantly improves detection accuracy and the quality of the rationales provided, compared to existing baselines. Additionally, we assess the generalizability of our framework across different LLM backbones, offering valuable insights for further optimization. Ovoerall, the key contributions of this work are summarized as follows:
  • We propose a novel framework, Lrq-Fact, to address the scalability issue in multimodal fact-checking by automatically generating multi-aspect questions.
  • Lrq-Fact incorporates a rule-based decision-maker that combines insights from both modalities, providing comprehensive rationales to examine overall veracity.
  • Extensive experiments on two benchmarks show improved detection accuracy and rationale quality compared to baselines. We also evaluate the generalizability of our framework across different LLM backbones.
@@ 2.1 Fact-Checking
Fact-checking is a critical task in combating misinformation, traditionally performed by human experts who generate questions and answers to verify the veracity of claims [18]. This process involves evaluating statements, comparing them with credible sources, and identifying inconsistencies [19]. However, manual fact-checking is labor-intensive and does not scale well with the growing online content [72]. Automated fact-checking systems have emerged to address this issue, with early approaches focusing primarily on textual claims [20].
Several works have explored various methods for automating fact-checking by leveraging machine learning and natural language processing (NLP) [50, 26, 31]. For example, one work expanded fact-checking capabilities by integrating context retrieval with large-scale databases to verify the factuality of statements [51].
However, while text-based fact-checking has made significant strides, existing approaches often fall short when dealing with multimodal content—such as the combination of images and text commonly found on social media platforms [79, 10]. This limitation has led to the need for advanced fact-checking systems that analyze both visual and textual elements.
2.2 Multimodal Misinformation
Misinformation has become a critical issue that affects various domains, from politics to public health [24, 38, 14]. Misinformation involves spreading false or misleading information, and can exist in different modalities—such as text, images, audio, and video—making it increasingly complex to detect [70, 60, 81].
Several datasets have been developed to evaluate text-based misinformation detection. For example, the FEVER dataset [68] consists of manipulated Wikipedia sentences annotated for veracity. Similarly, FakeNewsNet [63] collects misinformation from fact-checking websites and social media platforms. However, these datasets are limited to textual modality. Some recent datasets have focused on multimodal misinformation [23, 58]. The Fakeddit dataset [49] includes Photoshop-edited images from Reddit. DGM4 [61] introduces both global and local manipulations in images to alter sentiment and semantics, providing a comprehensive evaluation of multimodal misinformation. MMFakeBench [45] is designed to assess mixed-source multimodal misinformation. By including both synthetic and naturally occurring misinformation, MMFakeBench provides a valuable resource for evaluating modern VLMs and multimodal misinformation detection frameworks. These benchmarks highlight the need for detecting both textual and visual veracity distortions, as well as cross-modal consistency, which involves verifying if the visual and textual content align correctly.
In terms of detection, recent approaches aim to evaluate the consistency between different modalities. These techniques involve learning joint multimodal representations, where both visual and textual features are fused to detect inconsistencies or manipulations. Approaches like multimodal feature representation learning [1] and image captioning models [87] attempt to capture this alignment. However, these models often suffer from a lack of interpretability, making it difficult to interpret their decisions. To address this limitation, interpretable misinformation detection has gained attention [43, 15].
2.3 LLMs and VLMs
LLMs and VLMs have revolutionized the field of multimodal misinformation detection. LLMs, such as GPT-3 [7], GPT-4 [3], and LLaMA 3.1 [71], are designed to perform a wide range of language tasks, including summarization [16, 85], knowledge discovery [4, 32, 33, 66, 80, 82], question answering [69, 84] and various applications [83, 36, 12, 28]. These models are increasingly being used in fact-checking due to their ability to analyze textual claims. Recently, GPT-4o and LLaMA 3.1, in particular, have shown significant advancements in their ability to handle diverse, real-world misinformation by offering context-aware question generation and deeper understanding of semantic structures in texts [71, 5, 48, 3].
Meanwhile, VLMs, such as CLIP [57], ViLBERT [47], and Paligemma [6], have demonstrated their ability to bridge the gap between image and text by learning representations that align both modalities. These models are good at aligning fine-grained features between text and images, making them particularly useful for tasks such as verifying if an image’s content is consistent with its corresponding caption. By leveraging the capability of these VLMs in multimodal fact-checking, systems can detect subtle manipulations between visual and textual content [6].
In this work, we integrate LLMs and VLMs for multimodal misinformation detection. Recent studies have demonstrated the potential of leveraging their ability to generate diverse QA pairs based on both text and visual content [45, 64]. Inspired by this, Lrq-Fact combines LLMs like GPT-4o and LLaMA 3.1 for language understanding with VLMs like Paligemma for image analysis.
3 Methodology
In this section, we detail each module Lrq-Fact. As shown in Figure 2, Lrq-Fact integrates LLMs and VLMs to automatically generate relevant questions and answers (QAs) for both visual and textual content. It consists of four modules: (1) Image Description Generation, (2) Image-Focused QAs Generation, (3) Text-Focused QAs Generation, and (4) Rule-Based Decision-Maker. The prompt templates we designed for VLMs and LLMs are provided in Appendix (see Figure 8, 9, 10, 11, 12, and 13).
Figure 2: The overview pipeline of our Lrq-Fact framework consists of four key components: (a) Image Description, which provides detailed contextual descriptions of the image; (b) Image-Focused QAs, aimed at assessing the accuracy of the visual content; and (c) Text-Focused QAs, which detect textual inaccuracies, contradictions, or unsupported claims. Finally, all the gathered information is synthesized in (d) the Final Judgment & Explanation module, where a role-based decision-maker generates both the prediction results and comprehensive explanations.
@@ 3.1 Image Description Generation
In the first stage of our framework, we generate a detailed description of the image to assess consistency between the visual and textual content within our rule-based decision-maker module. This step is essential for identifying misinformation, as discrepancies between the image and the accompanying text are often a key source of misleading information. To achieve this, the VLM generates a comprehensive summary of the image, capturing key elements and describing the overall context of the scene depicted.
Formally, given the image $img_{i}$, we prompt the VLM to generate its corresponding description $des_{i}$:
(3.1) $des_{i}={\rm VLM}(img_{i}).$
3.2 Image-Focused QAs Generation
In the second stage of our framework, we focus on generating image-focused questions to verify the visual content related to the news article. Leveraging the capabilities of LLMs, we formulate queries targeting key aspects of the image. These questions are grounded in the context provided by the news article and are designed to verify the alignment of the image with the claims made in the text. We achieve this by employing in-context learning, where the LLM is provided with example questions that explore critical aspects of the image, such as objects, individuals, settings, and interactions. By including these examples in the prompt, the LLM gains a clearer understanding of the types of questions required, enabling it to generate new, relevant questions tailored to the context of the article  [5]. Intuitively, this approach ensures that the generated questions comprehensively address the accuracy, consistency, and authenticity of the image.
To generate appropriate questions, we also create a well-built prompt that guides the LLM to focus on extracting relevant information from the image. The prompt aims to instruct the LLM to create high-quality, targeted questions that help fact-checkers or automated systems determine whether the visual content supports, contradicts, or diverges from the claims in the news article. These questions probe specific details such as interactions between people, the setting, actions depicted, symbols, and the presence or absence of individuals or objects mentioned in the text. For instance, consider an article that describes a sporting event: “An umpire is in jump action while the pitcher launches the ball.” In this scenario, the LLM might generate the following questions to explore the visual evidence:
  • What sport is being played in the image?
  • Is the umpire in the image wearing the appropriate attire for the sport mentioned in the news article?
  • What is the action of the pitcher in the image?
  • Are there any visible spectators in the image?
  • Is the image likely to have been generated by AI, or is it a real photograph?
Intuitively, these questions help verify whether the image accurately represents the elements described in the article, ensure its details align with the expected aspects of the event, and examine the image for potential signs of manipulation, such as determining if it was AI-generated or shows signs of tampering.
Once the questions are generated, the VLM answers them by extracting critical visual details from the image, identifying key elements such as individuals, objects, the environment, and visual cues like the presence of spectators or the setting in response to the questions. This information helps assess the accuracy of the image content in relation to the article’s claims. The abovementioned QAs generation can be expressed as:
(3.2) ${ques_{i}^{v_{1}},...ques_{i}^{v_{m}}}={\rm LLM}(text_{i}),$
(3.3) $ans_{i}^{v_{j}}={\rm VLM}(img_{i},ques_{i}^{v_{j}}),\quad 0<j<m.$
3.3 Text-Focused QAs Generation
In the third stage of our framework, we focus on verifying the textual content of the news article by generating questions that challenge the factual claims presented in the text. Leveraging the capabilities of LLMs, these questions target specific details, such as dates, names, events, and locations, to critically assess the veracity of the article. This mimics the process human fact-checkers would follow, ensuring that the text is thoroughly examined for potential inaccuracies or unsupported claims. We achieve this by employing in-context learning, where the LLM is provided with example questions that explore critical aspects of the news article, such as factual claims, timelines, or significant events. For example, if the article describes a sporting event involving an umpire and a pitcher, the LLM will generate the following types of questions to examine the plausibility of these claims:
  • Is it common for an umpire to be in a “jump action” during a pitch, and does this align with standard umpiring practices in baseball?
  • Does the description of the pitcher “launching the ball” accurately reflect the typical terminology and actions used in baseball pitching?
  • Are there any known instances or rules in baseball that would require an umpire to jump while a pitch is being thrown?
These questions are designed to probe the accuracy of the article’s claims by evaluating them against known facts or commonly accepted standards. By analyzing the responses generated for these questions, the framework can identify potential discrepancies, factual errors, or unsupported statements, ensuring that the article does not spread misinformation.
Once the questions are generated, a single LLM is used to answer them based on its extensive knowledge:
(3.4) ${ques_{i}^{t_{1}},...ques_{i}^{t_{n}}}={\rm LLM}(text_{i}),$
(3.5) $ans_{i}^{t_{j}}={\rm LLM}(des_{i},ques_{i}^{t_{j}}),\quad 0<j<n.$
Intuitively, LLMs hold a vast repository of information, far beyond what any individual can retain, and provide access to this knowledge more intuitively and efficiently than traditional search engines [75]. This makes them highly effective for fact-checking, often supplying the extra factual knowledge needed to verify claims [50]. We hypothesize that LLMs, trained on large datasets, can serve as reliable substitutes for external evidence retrieval, performing fact-checking based solely on their parametric knowledge. This allows the LLM to provide relatively reliable answers to the generated questions, assessing the truthfulness of claims in the article and identifying textual inaccuracies, contradictions, or unsupported statements.

### v2

# Can LLMs Improve Multimodal Fact-Checking by Asking Relevant Questions?

@@ Abstract
Traditional fact-checking relies on humans to formulate relevant and targeted fact-checking questions (FCQs), search for evidence, and verify the factuality of claims. While Large Language Models (LLMs) have been commonly used to automate evidence retrieval and factuality verification at scale, their effectiveness for fact-checking is hindered by the absence of FCQ formulation. To bridge this gap, we seek to answer two research questions: (1) Can LLMs generate relevant FCQs? (2) Can LLM-generated FCQs improve multimodal fact-checking? We therefore introduce a framework Lrq-Fact for using LLMs to generate relevant FCQs to facilitate evidence retrieval and enhance fact-checking by probing information across multiple modalities. Through extensive experiments, we verify if Lrq-Fact can generate relevant FCQs of different types and if Lrq-Fact can consistently outperform baseline methods in multimodal fact-checking. Further analysis illustrates how each component in Lrq-Fact works toward improving the fact-checking performance.
Can LLMs Improve Multimodal Fact-Checking by Asking Relevant Questions?
Alimohammad Beigi1 Bohan Jiang1 Dawei Li1 Zhen Tan1 Pouya Shaeri1Tharindu Kumarage1Amrita Bhattacharjee1Huan Liu 1 1School of Computing, and Augmented Intelligence, Arizona State University {abeigi, bjiang14, daweili5, ztan36, pshaeri, kskumara, abhatt43, huanliu}@asu.edu
@@ 1 Introduction
Fact-checking is an important yet challenging task in combating online misinformation. Modern misinformation often spreads across multiple modalities, containing both textual and visual falsehoods, which significantly complicates accurate and efficient fact-checking Akhtar et al. (2023). In journalism, fact-checking is traditionally a three-step process Graves and Amazeen (2019), where human fact-checkers (1) formulate relevant and targeted fact-checking questions (FCQs), (2) search for supporting evidence, and (3) verify the factuality of claims or statements. Fact-checkers leverage domain knowledge to pose precise and contextually relevant FCQs, ensuring claims are evaluated from multiple perspectives PolitiFact.com (2011); Vlachos and Riedel (2014); Vo and Lee (2019). However, given the rapid proliferation of online misinformation Chen and Shu (2023); Jiang et al. (2024b), manual fact-checking is insufficient to keep pace with the scale of the problem Shaeri and Katanforoush (2023).
To improve efficiency, researchers have developed automated fact-checking (AFC) systems capable of identifying misinformation Hassan et al. (2017); Miranda et al. (2019); Dierickx et al. (2023). Recently, Large Language Models (LLMs) have been explored for zero-shot fact-checking Geng et al. (2024). However, research pointed out that directly prompting LLMs for fact-checking remains less effective in many cases Yao et al. (2023). One key issue is the absence of relevant FCQs, which are essential to guide LLMs in retrieving accurate supporting evidence and conducting reliable veracity evaluations Chen et al. (2022); Pan et al. (2023); Setty (2024). In this work, we investigate the potential of LLM-generated FCQs by answering two research questions (Figure 1):
  • Are LLMs capable of generating relevant FCQs?
  • Can the generated FCQs improve AFC systems?
Figure 1: The two research questions we aim to address in this work.
Inspired by the human fact-checking process, we introduce LLM-generated Relevant Questions for multimodal FACT-checking (Lrq-Fact), an LLM-based framework designed to automatically generate relevant and targeted FCQs to guide the AFC system to fact-check multimodal misinformation. Lrq-Fact first generates two types of FCQs: (1) visual FCQs, which assess whether an image accurately represents critical details such as people, objects, or events mentioned in the text, and (2) textual FCQs, which question whether the textual claims or statements are supported by evidence. Human annotators and LLM judges evaluate the quality of LLM-generated FCQs with pre-defined rules, and show that most of the textual and visual FCQs generated by Lrq-Fact are contextually relevant to the fact-checking task.
Next, we seek to answer whether the generated questions can improve multimodal fact-checking. With the guidance of relevant FCQs, Lrq-Fact incorporates the internal training knowledge of LLM and external online searching to strengthen its evidence retrieval and verification capabilities. The up-to-date online information is particularly valuable when fact-checking claims related to emerging or rapidly evolving events, where LLMs often lack sufficient ground truth knowledge. Extensive experiments are conducted on three datasets. Our results show that Lrq-Fact can outperform baseline methods significantly. Furthermore, our ablation study probes into the model’s modular components to evaluate their effectiveness in fact-checking performance. Last but not least, we demonstrate that Lrq-Fact is highly adaptive, generalizing across different LLM backbones. In summary, the key contributions of this work are as follows:
  • We analyze the use of LLMs to generate relevant and targeted FCQs.
  • We explore and experiment the effectiveness of LLM-generated FCQs in multimodal fact-checking on three benchmark datasets.
  • Further analysis illustrates how each component in Lrq-Fact contributes to performance.
@@ 2 Related Work
2.1 Multimodal Misinformation
Misinformation spans multiple domains, consisting of various modalities such as text and images, making detection increasingly complex Li et al. (2020); Jiang et al. (2024a); Tufchi et al. (2023). While early misinformation detection mainly focused on textural content Thorne et al. (2018); Shu et al. (2020), recent datasets have incorporated multimodal misinformation, such as Fakeddit Nakamura et al. (2019), DGM4 Shao et al. (2023), and MMFakeBench Liu et al. (2024b). Multimodal misinformation detection methods have been developed to learn joint representations of different modalities Abdali et al. (2022); Zhou et al. (2020). Some studies proposed explainable detection frameworks to enhance the interpretability Liu et al. (2023); Fung et al. (2021). However, these models may fall short when facing newly emerged or rapidly evolving topics. Our method addresses this issue by utilizing up-to-date Google Search as an external knowledge source.
@@ 2.2 Fact-Checking
Fact-checking is essential for combating misinformation, traditionally relying on human fact-checkers to verify claims by generating FCQs and cross-referencing credible sources Graves (2018); Graves and Amazeen (2019). However, manual fact-checking is resource-intensive and struggles to scale with increasing online misinformation Vlachos and Riedel (2014). AFC systems address this challenge, with early approaches focusing on textual claims using Machine Learning and Natural Language Processing techniques Guo et al. (2022); Nakov et al. (2021); Karadzhov et al. (2017). Recent studies have expanded AFC’s capabilities by integrating large-scale evidence retrieval Nie et al. (2019); Akhtar et al. (2023); Geng et al. (2024). However, the lack of FCQs limited the performance of AFC. We address this limitation by harnessing LLMs to generate relevant and targeted FCQs.
2.3 Language Models for Fact-Checking
LLMs and Vision Language Models (VLMs) have had significant impacts on AFC. LLMs from the GPT family Brown (2020); Achiam et al. (2023); Hurst et al. (2024) and LLaMA series Touvron et al. (2023); Vavekanand and Sam (2024) excel in language understanding and context-aware question generation, enhancing AFC performance Achiam et al. (2023); Vavekanand and Sam (2024); Beigi et al. (2024). VLMs such as CLIP Radford et al. (2021), ViLBERT Lu et al. (2019), and Paligemma Beyer et al. (2024), enable cross-modal analysis, aligning visual and textual features to detect false statements. Recent studies highlight the potential of combining LLMs and VLMs to generate targeted FCQs for better fact-checking Singh et al. (2021); Chen et al. (2022); Pan et al. (2023). Our framework utilizes LLMs and VLMs to generate relevant FCQs, improving multi-modal fact-checking.
Figure 2: Human and GPT-4o Question Quality Evaluations Across Datasets (50 Samples per Dataset-Modality).
3 Task Definition
We define the task of multimodal fact-checking as a multiclass classification problem. Given a news article $text_{i}$ and accompanying image $img_{i}$ as input, we aim to classify the news into one of the following categories:
  • Real ($y=0$): The news is factually accurate and consistent.
  • Textual Veracity Distortion (TVD, $y=1$): False or misleading claims in the text.
  • Visual Veracity Distortion (VVD, $y=2$): Manipulated or misleading images.
  • Cross-Modal Mismatch (CMM, $y=3$): Inconsistencies between text and image.
This approach allows for a finer classification of misinformation, improving targeted fact-checking.
4 RQ1: Can LLMs Generate Relevant FCQs?
In this section, we discuss the FCQ generation phase of Lrq-Fact. Specifically, Lrq-Fact produces visual (Sec. 4.1) and textual FCQs (Sec. 4.2). We then evaluate the quality of the generated FCQs using a combination of LLM-based and human assessments (Sec. 4.3).
4.1 Visual FCQs Generation
In the first stage, Lrq-Fact formulates targeted visual FCQs based on the news article to verify whether the visual content aligns with claims in the news article. Importantly, the LLM does not have direct access to the image itself; rather, it generates questions based solely on the article’s content, anticipating what aspects might be depicted in an accompanying image. To guide this process, we employ a structured prompt (see Figure 12 in Appendix A.2) that instruct the LLM to focus on objects, settings, interactions, and potential manipulations, to enable a comprehensive verification of accuracy, consistency, and authenticity.
For instance, given a news describing a sporting event, the LLM may generate questions such as:
  • What sport is being played in the image?
  • Is the pitcher actively throwing the ball?
  • Are there visible spectators?
  • Is the image AI-generated or a real photograph?
4.2 Textual FCQs Generation
The second stage generates fact-checking questions that examine the factual accuracy of textual claims, including dates, names, and events, mimicking human verification methods. Leveraging in-context learning, the LLM formulates precise queries to validate these claims, guided by a structured prompt that ensures relevance and depth in question generation (see Figure 14 in Appendix A.2). For example, if an article states that an umpire performed a “jump action” during a pitch, the LLM might ask:
  • Is it common for an umpire to jump during a pitch in baseball?
  • Does “launching the ball” align with standard baseball terminology?
  • Are there baseball rules requiring an umpire to jump while a pitch is thrown?
4.3 FCQ Quality Evaluation
To ensure a cost-efficient evaluation process, we use human annotators and LLM-as-a-judge to assess the relevance of the generated FCQs. We first establish structured evaluation criteria by incorporating best practices from leading fact-checking organizations Snopes (2024); PolitiFact (2024); FactCheck (2024). More details can be found in Table 4 in Appendix A.4. Specifically, we employ ten predefined criteria to guide the LLM judge in FCQ quality evaluation. These criteria cover aspects such as critical thinking, analytical depth, precision, factual accuracy, logical consistency, and source credibility (see Figure 17 in Appendix A.2).
To validate the reliability of the LLM judge, we compare the results of LLM and human annotators (Figure 2) using Fleiss’ Kappa correlation Fleiss (1971). In particular, we randomly sample 50 instances from each dataset, with each instance containing five visual and five textual FCQs. Two human annotators evaluate these FCQs to decide whether they satisfy all predefined criteria.
QuestionsDataset MMFakeBench DGM4 Factify
Textual 0.78 0.83 0.80
Visual 0.79 0.82 0.81
Table 1: Fleiss’ Kappa score between human and LLM evaluations of question relevancy.
Our findings show that LLM-generated FCQs are highly relevant (Table 1), with substantial agreement between human annotators and LLM judge. The high Fleiss’ Kappa scores show the potential to extend the LLM-based quality evaluation to all FCQs. As shown in Figure 3, nearly $73\%$ of visual and $93.6\%$ of textual FCQs are relevant. These results highlight the effectiveness of LLMs in generating high-quality FCQs.
Figure 3: GPT-4o Evaluation of Question Relevance Across Datasets (1000 Samples per Dataset-Modality).
5 RQ2: Can LLM-Generated FCQs improve Multimodal Fact-Checking?
In this section, we first present the remaining components of Lrq-Fact, which includes answering the generated FCQs (Sec. 5.1, 5.2, and 5.3) and a Rule-Based Decision-Maker (Sec. 5.4). We then investigate RQ2 with extensive experiments and analysis on three datasets. We also provide a case study to showcase the effectiveness of FCQs (Sec. 5.8).
@@ 5.1 Image Description Generation
Since our framework performs the fact-checking in the textual space, we first generate a detailed textual description of the image. The aim is to identify the scene/content/action in the image so that the rule-based stage of the pipeline can use this information to check for consistency between information depicted in the news text vs. shown in the news image. To achieve this, we prompt (see Figure 11 in Appendix A.2) the VLM to generate a summary of the image, instructing it to ensure it to capture key elements of the scene depicted. Formally, given the image $img_{i}$, we use the VLM to generate its corresponding description $des_{i}$:
$des_{i}={\rm VLM}(img_{i}).$ (1)
Figure 4: The proposed framework, Lrq-Fact, draws on insights from human fact-checking process.
5.2 Answering Visual FCQs via VLM
To answer the Visual FCQs, a VLM extracts relevant visual details, allowing for a clear evaluation of how well the image matches the text:
${ques_{i}^{v_{1}},...ques_{i}^{v_{m}}}={\rm LLM}(text_{i}),$ (2)
[…]
5.3 Answering Textual FCQs via RAG
While LLMs are powerful in language generation, they are prone to hallucinations, leading to inaccurate or unverified information. To enhance factual reliability, we employ Retrieval-Augmented Generation (RAG), which grounds responses in external, verifiable sources. This is especially important for newly emerging topics in news articles since the LLM may not have knowledge of such topics due to its knowledge cutoff. We implement an automated online search by using the Google Web Search API Google to gather relevant news articles for each claim. We identify the top 10 sources, extract their textual content, and compile a factual document containing relevant information. Next, we use LangChain Chase (2022) to retrieve the most relevant passages from this document for fact-checking questions, ensuring that only contextually accurate information is used for verification. Finally, an LLM generates answers based on the retrieved content, reducing hallucinations and improving factual accuracy. To guide the LLM effectively, we employ a carefully designed prompt (see Figure 16 in Appendix A.2) that instructs the model to rely solely on the provided factual evidence. This process is formulated as follows:
$Doc_{i}={\rm SearchOnline}(text_{i})$ (4)
$RelContent_{i}$ $={\rm Retriever}(Doc_{i},ques_{i}^{t_{j}})$ (5)
$ans_{i}^{t_{j}}$ $={\rm LLM}(RelContent_{i},ques_{i}^{t_{j}})$
$\quad\quad\quad 0<j<n.$